Calculate the x-component and the y-component (in m) of the vector with magnitude 10.0 m and direction 30.0°.


1
Expert's answer
2021-09-21T06:33:25-0400
"r_x=rcos\theta=10.0\ m\cdot cos30^{\circ}=8.66\ m,""r_y=rsin\theta=10.0\ m\cdot sin30^{\circ}=5.0\ m."